OpenStudy (anonymous):

write an equation that is parallel to 3x-y=-6 and contains the points (0,0)

OpenStudy (anonymous):

ok

OpenStudy (gorv):

welcome to os :)

OpenStudy (gorv):

parallel line have same slope

OpenStudy (gorv):

so bring it in standard form of y=mx+c

OpenStudy (gorv):

add y on both side 3x=-6+y add 6 both side

OpenStudy (gorv):

y=3x+6 so slope =3

OpenStudy (gorv):

so our parallel line y=3x+c

OpenStudy (anonymous):

thank you. I need an equation that is parallel and contains the points (0,0)

OpenStudy (texaschic101):

y = mx + b slope(m) = 3 (0,0)...x = 0 and y = 0 now sub y = mx + b 0 = 3(0) + b 0 = b so your parallel equation is : y = 3x + 0 or just y = 3x